OpenStudy (anonymous):

When x->0 lim TAN(x2)/ x =

OpenStudy (amistre64):

tan = sin/cos sin/cos sin ------ = ----- x x cos sin(x^2) = (x^2) - (x^2)^3/3! + (x^2)^5/5! - ... x cos(x^2) = x - x(x^2)^2/2! + x(x^2)^4/4! + x(x^2)^6/6! - ... i spose we can try a Lhopital?

OpenStudy (amistre64):

tan(x^2) to sec^2(x^2) 2x x to 1 looks to go zero to me

OpenStudy (amistre64):

we could have gone with the series of tan(x^2)/2 and seen that its a poly of x terms